# Quantum enhanced real-time sensing of protein binding process

### v1

@@ Acknowledgements.
Appendix
The numerical simulation in Fig 1b is based on the following theoretical framework.
@@
The characteristic matrix $M^{\prime}$ of the stratified medium here gold film bounded by two semi-infinite media on both sides as shown in Fig 9 evaluated at the gold-prism interface is given by,
@@
Here, $\epsilon_{l}$ is the dielectric permittivity of the medium l=1,2, $h$ is the thickness of the gold film and $\omega$ is the frequency of the incident field. In our calculations, the magnetic permeability $\mu_{l}$ is considered to be 1 since the magnetic field is assumed to be not interacting in this phenomenon.
The reflection coefficient r at the interface between gold and prism is given by the equation,
@@
The reflectivity which is the $\mid r\mid^{2}$ is then plotted against $\theta_{1}$. This quantity reaches its minimum at the resonance angle where the energy from the incident electric field is mostly absorbed by the excited surface plasmon polaritons. While calculating the SPR angle we also consider the Fresnel coefficients at the prism-air interface at which the EM field enters and exits the prism. Thus we see the reflectivity for the deionized water reaches its minimum at the angle of 66 degrees and shifts towards right after the BSA binds the gold.

### v2

@@ Acknowledgements.
Appendix A: Transfer-Matrix method
The Kretschmann geometry of the SPR sensor can be considered as a multilayer system where gold film with refractive index $n_{2}$ is situated between two semi-infinite media namely the prism with refractive of glass $n_{1}$ and the sample for example deionized water with refractive index of $n_{3}$ as shown in Fig. 10. The reflectance and transmission of this multilayer system can be calculated using the Transfer-Matrix method (TMM) [35]. TMM is a systematic way of finding the solutions for the electromagnetic wave for appropriate boundary conditions at each interface in the stratified medium. The stratified medium is characterized by a matrix $M^{\prime}$ called as the characteristic matrix of the system. The reflectance curve as a function of incident angle shown in Fig. 1(b) is plotted using this method.
@@
Here, the characteristic matrix $M^{\prime}$ of the stratified medium evaluated at the gold-prism interface is given by,
@@
Here, $\epsilon_{l}$ is the dielectric permittivity of the medium $l=1,2,3$, $h$ is the thickness of the gold film and $\omega$ is the frequency of the incident field. In our calculations, the magnetic permeability $\mu_{l}$ is considered to be 1 since the magnetic field is assumed to be not interacting in this phenomenon. $\theta_{l}$ with $l=1,2,3$ are the angles made by the EM field with the normal for the respective mediums as shown in Fig. 10.
Using TMM the reflection coefficient r at the interface between gold and prism is given by the equation,
@@
The reflectivity which is the $|r|^{2}$ is then plotted against $\theta_{1}$. This quantity reaches its minimum at the resonance angle where the energy from the incident electric field is mostly absorbed by the excited surface plasmon polaritons. While calculating the SPR angle we also consider the Fresnel coefficients at the prism-air interface at which the EM field enters and exits the prism. Thus we see the reflectivity in Fig. 1(b) for the deionized water reaches its minimum at the angle of 66.73 degrees and shifts towards right after the BSA binds the gold.
Appendix B: Sensor assembly
We prepared the sensor using the BK7 (n=1.51) right angle prism from Thorlabs, a microscope glass slide with 50 nm gold layer (SKU: AU.0500.ALSI) bought from Platypus technologies, and the customized flow cell designed and fabricated using on campus 3D printing facility. As shown in Fig. 3(c), the sample delivery system needs to work such that the proteins get in contact with the gold without leaking from the flow cell. We use O-ring to avoid any leakage from the flow cell as it is tightly clamped against the microscope slide and the prism. The input and the output ports of the flow cell are connected to a peristaltic pump to control the flow of the sample. Another challenge during the clamping process is the formation of air bubbles between the glass slide and the prism. We use immersion oil (56822-50ml Sigma Aldrich, n=1.5) to avoid any air bubbles. We chose the refractive index of the glass portion of the microscope slide same as the refractive index of prism as well as the immersion oil such that there is no refraction at the gold-dielectric interface.
Figure 11: The shot noise limit (dB) associated with the coherent source was plotted against the total intensity (V) of the source entering the detectors A and B.
Appendix C: Shot Noise Limitation
We verified the coherent source for its shot-noise limit. The two coherent beams were directed in the two detectors A and B replicating the arrangement for TMBSS source. The verification was done using the oscilloscope for voltage readings and the spectrum analyzer was used to record the noise power around the working frequency of 2MHz. As shown in Fig. 11 the noise power plotted on y axis is taken relative to the noise power corresponding to the total intensity of 1.5 V. As we can see the SNL increases by 3dB when the intensity is doubled demonstrating the linear dependence of the variance with respect to the mean of the signal.
Appendix D: SPR characteristics for different concentrations
Figure 12: Reflectivity of the sensor is plotted against the relative scanning angle demonstrating the shift in the resonance angle from $0\%$ BSA (deionized water) to $10\%$ BSA solution.
We recorded the reflectivity of the sensor against the scanning angle for $0\%$ BSA (deionized water) and $10\%$ BSA solution as shown in Fig. 12 using the coherent source. Next, we provide the theoretical expectation for the behavior of the SPR resonance angle for various concentrations of BSA solution as explained below.
We use the information provided in [42] to find the refractive index of various concentrations of the BSA solution as shown in Fig. 13(a). We then map these refractive indices to the angle of resonance calculated using the Transfer Matrix Method (TMM) as explained in appendix A and the schematics section in this paper. Thus we plot the resonance angle as a function of the concentration of BSA in water as shown in Fig. 13(b). The plot shows the linearity as expected for a calibration curve.
Figure 13: (a) Refractive index for different concentrations of BSA solution was plotted based on the findings reported in [42]. (b) The expected angle of resonance was plotted for 10 different concentrations of BSA solution based on the transfer matrix method applied to the geometry of SPR sensor.